OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim? If D = dad's age now and J = Jim's age now, which of the following systems could be used to solve the problem?

OpenStudy (anonymous):

Jim is 15 years old

OpenStudy (anonymous):

j * 4 = d j + 10 = d2 + 20